# Multi-Agent Reinforcement Learning from Human Feedback: Data Coverage and Algorithmic Techniques

### v2

@@ Abstract
We initiate the study of Multi-Agent Reinforcement Learning from Human Feedback (MARLHF), exploring both theoretical foundations and empirical validations. We define the task as identifying Nash equilibrium from a preference-only offline dataset in general-sum games, a problem marked by the challenge of sparse feedback signals. Our theory establishes the upper complexity bounds for Nash Equilibrium in effective MARLHF, demonstrating that single-policy coverage is inadequate and highlighting the importance of unilateral dataset coverage. These theoretical insights are verified through comprehensive experiments. To enhance the practical performance, we further introduce two algorithmic techniques. (1) We propose a Mean Squared Error (MSE) regularization along the time axis to achieve a more uniform reward distribution and improve reward learning outcomes. (2) We utilize imitation learning to approximate the reference policy, ensuring stability and effectiveness in training. Our findings underscore the multifaceted approach required for MARLHF, paving the way for effective preference-based multi-agent systems.
@@ 1 Introduction
Despite some empirical successes in agent systems utilizing closed-source LLMs, finetuning these systems and aligning them with human preferences remains a challenge. Reinforcement learning from human feedback (RLHF) has played an important role in aligning LLMs with human preferences (Christiano et al., 2017; Ziegler et al., 2019). However, unexpected behavior can arise when multiple LLMs interact with each other. In addition, reward design has been a hard problem in multi-agent reinforcement learning (Devlin et al., 2011). Thus, it is crucial to further align the multi-agent system with human feedback.
We address this problem through both theoretical analysis and empirical experiments. Theoretically, we characterize the dataset coverage condition for MARLHF that enables learning the Nash equilibrium, which serves as a favorable policy for each player. Empirically, we validate our theoretical insights through comprehensive experiments utilizing the proposed algorithmic techniques.
@@ 1.1 Contributions and Technical Novelties
1. Necessary and Sufficient Dataset Coverage Condition for MARLHF.
In single-agent RLHF, (Zhu et al., 2023) demonstrated that single policy coverage is sufficient for learning the optimal policy. However, we prove that this condition no longer holds for MARLHF by providing a counterexample. Instead, we introduce an algorithm that operates under unilateral coverage, a condition derived from offline MARL (Cui and Du, 2022a; Zhong et al., 2022). Specifically, this condition requires the dataset to cover all unilateral deviations from a Nash equilibrium policy. For further details, see Section 4.
Figure 1: The overall pipeline of offline MARLHF. $\mathcal{D}$ is the preference dataset where $\tau$ is a one trajectory and $\mathbf{y}_{i}\in\{1,-1\}^{m}$ indicates which trajectory is preferred by each agent. $r_{\phi}$ is the learned reward. $\pi_{ref}$ is the learned reference policy using imitation learning.
@@
As a foundational exploration into MARLHF research, we focus on employing the simplest learning framework, incorporating only the essential techniques necessary to ensure the approach’s feasibility. The framework consists of three key components: 1) leveraging the preference dataset to learn a reward function, 2) mitigating extrapolation errors using a reference policy, and 3) determining the final policy. Figure 1 provides an overview of the process.
However, additional algorithmic techniques are needed in order to find a strong policy even if we have a dataset with good coverage according to the theory.
@@
  • Imitation learning policy as a reference. We choose to follow the standard RLHF work (Wang et al., 2023b), where the extrapolation problem was alleviated by adding an extra KL reward to the final optimization objective. Since we do not have access to the actual reference model, we implement imitation learning across the entire dataset to approximate it. The final policy is trained using a DQN-based Value Decomposition Network (VDN) (Mnih et al., 2013; Sunehag et al., 2017). Our ablation study demonstrates the effectiveness of selecting an appropriate KL reward coefficient (see Table 3).
@@
Our experiments, following the pipeline described above, confirm the theoretical necessity of unilateral coverage. We conducted comprehensive ablation studies on three cooperative Multi-Agent Particle Environment (MPE) scenarios (Mordatch and Abbeel, 2017): Spread-v3, Tag-v3, and Reference-v3. These studies focused on the hyperparameter selection for the reward regularization coefficient, KL reward coefficient, and dataset diversity. The empirical results (Table 2) demonstrate that: 1) simply adding trivial trajectories to expert demonstrations can enhance performance, 2) unilateral datasets are advantageous, and 3) dataset diversity contributes to lower variance.
Our ablation experiments underscore the effectiveness of the proposed algorithmic techniques. Additionally, we tested the sensitivity of the hyperparameters used in MSE and KL regularization, and we introduce a principled standardization technique that can efficiently tune hyperparameters across all environments and datasets.
@@ 2 Related Works
RLHF (or preference-based RL) plays a pivotal role in alignment with various tasks such as video games (Warnell et al., 2018; Brown et al., 2019), robotics (Jain et al., 2013; Kupcsik et al., 2016; Christiano et al., 2023; Shin et al., 2023), image augmentation (Metcalf et al., 2024), and large language models (Ziegler et al., 2020; Wu et al., 2021; Nakano et al., 2022; Menick et al., 2022; Stiennon et al., 2022; Bai et al., 2022; Glaese et al., 2022; Ganguli et al., 2022; Ouyang et al., 2022). Additionally, a body of work focuses on the reward models behind preference data (Sadigh et al., 2017; Bıyık and Sadigh, 2018; Gao et al., 2022; Hejna and Sadigh, 2023). Recently, direct preference optimization (DPO, Rafailov et al. (2023)) and its variants (Azar et al., 2023; Rafailov et al., 2024) approach RLHF without directly handling the reward model. Theoretical studies have also explored guarantees, such as sample complexity and regret, and the limitations of certain RLHF algorithms (Novoseller et al., 2020; Xu et al., 2020; Pacchiano et al., 2023; Chen et al., 2022; Razin et al., 2023; Zhu et al., 2024a; Wang et al., 2023b; Xiong et al., 2024; Zhu et al., 2024b).
@@
Offline RL (Lange et al., 2012; Levine et al., 2020) has achieved success in a wide range of real-world applications, including robotics (Pinto and Gupta, 2015; Levine et al., 2016; Chebotar et al., 2021; Kumar et al., 2023), healthcare (Raghu et al., 2017; Wang et al., 2018), and autonomous driving (Shi et al., 2021; Lee et al., 2024). Key algorithms such as Behavior Cloning, BRAC (Wu et al., 2019), BEAR (Kumar et al., 2019), and CQL (Kumar et al., 2020; Lyu et al., 2024) have driven these successes. Theoretical research on offline RL has primarily focused on sample complexity under various dataset coverage assumptions Le et al. (2019); Chen and Jiang (2019); Yin et al. (2020); Rashidinejad et al. (2023); Yin et al. (2021, 2022); Shi et al. (2022); Nguyen-Tang et al. (2022); Xie et al. (2022); Xiong et al. (2023); Li et al. (2024).
@@
Many real-world scenarios are naturally modeled as multi-agent environments, whether cooperative or competitive. As a result, MARL has gained popularity in video games (Tian et al., 2017; Vinyals et al., 2017; Silver et al., 2017; Vinyals et al., 2019), network design (Shamsoshoara et al., 2018; Kaur and Kumar, 2020), energy sharing (Prasad and Dusparic, 2018), and autonomous driving (Palanisamy, 2019; Yu et al., 2020; Zhou et al., 2022). Prominent algorithms in MARL include IQL (Tan, 2003), MADDPG (Lowe et al., 2020), COMA (Foerster et al., 2017), MAPPO (Yu et al., 2022), VDN (Sunehag et al., 2017), and QMIX (Rashid et al., 2018).
@@
Offline MARL is a practical solution for handling sophisticated multi-agent environments. Empirically, to address issues related to out-of-distribution actions and complex reward functions, previous works have developed algorithms such as MABCQ (Jiang and Lu, 2023), ICQ-MA (Yang et al., 2021), OMAR (Pan et al., 2022), and OMIGA (Wang et al., 2023a), which incorporate regularization or constraints on these actions and functions. MOMA-PPO (Barde et al., 2024) is a model-based approach to offline MARL that generates synthetic interaction data from offline datasets. Tseng et al. (2022) combines knowledge distillation with multi-agent decision transformers (Meng et al., 2022) for offline MARL. Theoretical understanding of offline MARL, particularly in the context of Markov games, has been advanced by works that provide sample complexity guarantees for learning equilibria Sidford et al. (2019); Cui and Yang (2020); Zhang et al. (2023a, 2020); Abe and Kaneko (2020); Cui and Du (2022a, b); Zhang et al. (2023b); Blanchet et al. (2023); Shi et al. (2023).
@@ 4 Dataset Coverage Theory for MARLHF
In this section, we study the dataset coverage assumptions for offline MARLHF. For offline single-agent RLHF, (Zhu et al., 2023; Zhan et al., 2023) shows that single policy coverage is sufficient for learning the optimal policy. However, we prove that this assumption is insufficient in the multi-agent setting by constructing an counterexample. In addition, we prove that unilateral policy coverage is adequate for learning the Nash equilibrium.
@@ 5 Algorithmic Techniques for Practical Performance
In Section 4, we provided a theoretical characterization of the dataset requirements for MARLHF. However, the algorithm used in Theorem 2 is not computationally efficient. In this section, we propose a practical algorithm for MARLHF and validate our theoretical findings through experiments.
@@ 5.1 High-level Methodology
Our MARLHF pipeline consists of two phases: In the first step, we train a reward prediction model $\phi$ and approximate the behavior policy $\pi^{b}$ using imitation learning; in the second step, we then apply an MARL algorithm to maximize a combination of the KL-divergence-based reward and standardized predicted reward $r_{\phi}$, ultimately deriving the final policy $\pi_{\mathbf{w}}$.
Step 1: Reward Training and Reference Approximation. Given the preference signals of trajectories, we use neural networks to predict step-wise rewards $r_{\phi}(s_{h},a_{h})$ for each agent, minimizing the loss defined in (1). The objective is to map $(s,a_{i})$-pairs to reward values such that the team returns align with the preference signals. At the same time, in order to utilize KL reward $-\log\frac{\pi_{\mathbf{w}}(s,a)}{\pi^{b}(s,a)}$ to cope with the extrapolation error in offline learning, an imitation learner $\pi_{\text{ref}}$ is trained over the entire dataset to model the behavior policy $\pi^{b}$.
Step 2: Offline MARL. Although in this work, VDN is chosen as the MARL oracle, it should be noted that other MARL architectures are also applicable. With the reward model $r_{\phi}$ and the approximated reference policy $\pi_{\text{ref}}$ learned in the first step, we are now able to construct a virtual step-wise reward for each agent. The agents are then trained to maximize the target defined in (3).
@@ 5.2 Reward Regularization
5.3 Imitation Learning Policy as Reference
There are various methods to mitigate the over-extrapolation errors in offline RL (Peng et al., 2019; Nair et al., 2021), including conservative loss over the Q-function (Kumar et al., 2020) and directly restricting the learned policy actions to those within within the dataset (Fujimoto et al., 2019). For simplicity and consistency with the former RLHF framework (Ouyang et al., 2022), we adopt the same per-step KL reward for enforcing restrictions between $\pi^{b}$ and $\pi_{\mathbf{w}}$. In many scenarios, a direct reference to the behavior policy is not accessible. As an alternative, imitation learning is used to estimate the reference policy, $\pi_{\text{ref}}\approx\pi^{b}$. To stabilize training, we standardize $r_{\phi}$ over $\mathcal{D}$ before combining it with the KL reward to make it comparable:
$\text{objective}(\mathbf{w})=\mathbb{E}_{\tau\sim\pi_{\mathbf{w}}}\left[\sum_{h=1}^{H}r_{\text{std}}(s_{h},\mathbf{a}_{h},\phi)-\beta\log\frac{\pi_{\mathbf{w}}(s_{h},\mathbf{a}_{h})}{\pi_{\text{ref}}(s_{h},\mathbf{a}_{h})}\right],$ (3)
where $\beta$ is the KL reward coefficient, set to be (1, 1, 3) in Spread-v3, Reference-3 and Tag-v3 respectively, and the standardized reward $r_{\text{std}}$ is defined as:
@@ 5.3 Imitation Learning Policy as Reference
Notice that the explicit expression of $\pi_{\mathbf{w}}(s_{h},\mathbf{a}_{h})$ only exists in methods with an actor model, which leads to certain planning-based algorithms, such as Q-learning methods, being unable to directly use Equation 3. To enhance the compatibility of our pipeline, we substitute $1/|A|$ for $\pi_{\mathbf{w}}(s_{h},\mathbf{a}_{h})$ in such cases. Although this substitution formally removes the KL reward, in algorithms like Q-learning that output deterministic policies, the replaced reward is equivalent to encouraging the agent to match its deterministic choices with the actions most preferred by the reference policy. In other words, it seeks the closet element in the deterministic policy family to the reference policy, which is a degradation of the KL reward restricted in the domain of deterministic policy. The effectiveness of this method is validated in the ablation study (cf. Section 6.4).
@@ 6 Experiments
We design a series of experiments to validate our theories and methods in common general-sum games. Specifically, we first use VDN Sunehag et al. (2017) to train expert agents, and take intermediate checkpoints as rookie agents. Then, we use these agents to collect datasets and use the Broadley-Terry model over standardized returns to simulate human preference. Experiments are carried out to verify the efficiency of our approach with unilateral policy dataset coverage (in Theorem 2) while single policy coverage is insufficient (stated in Theorem 1). We also design ablation studies to showcase the importance of our methods, particularly focusing on reward structure regularization and imitation poclies as reference models.

### v3

# Preference-Based Multi-Agent Reinforcement Learning: Data Coverage and Algorithmic Techniques

@@ Abstract
We initiate the study of Preference-Based Multi-Agent Reinforcement Learning (PbMARL), exploring both theoretical foundations and empirical validations. We define the task as identifying the Nash equilibrium from a preference-only offline dataset in general-sum games, a problem marked by the challenge of sparse feedback signals. Our theory establishes the upper complexity bounds for Nash Equilibrium in effective PbMARL, demonstrating that single-policy coverage is inadequate and highlighting the importance of unilateral dataset coverage. These theoretical insights are verified through comprehensive experiments. To enhance the practical performance, we further introduce two algorithmic techniques. (1) We propose a Mean Squared Error (MSE) regularization along the time axis to achieve a more uniform reward distribution and improve reward learning outcomes. (2) We propose an additional penalty based on the distribution of the dataset to incorporate pessimism, improving stability and effectiveness during training. Our findings underscore the multifaceted approach required for PbMARL, paving the way for effective preference-based multi-agent systems.
@@ 1 Introduction
Despite some empirical successes in agent systems utilizing closed-source LLMs, finetuning these systems and aligning them with human preferences remains a challenge. Reinforcement learning from human feedback (RLHF) has played an important role in aligning LLMs with human preferences (Christiano et al., 2017; Ziegler et al., 2019). However, unexpected behavior can arise when multiple LLMs interact with each other. In addition, reward design has been a hard problem in multi-agent reinforcement learning (Devlin et al., 2011). Thus, it is crucial to further align the multi-agent system from preference feedback.
We address this problem through both theoretical analysis and empirical experiments. Theoretically, we characterize the dataset coverage condition for PbMARL that enables learning the Nash equilibrium, which serves as a favorable policy for each player. Empirically, we validate our theoretical insights through comprehensive experiments utilizing the proposed algorithmic techniques.
@@ 1.1 Contributions and Technical Novelties
1. Necessary and Sufficient Dataset Coverage Condition for PbMARL.
In single-agent RLHF, (Zhu et al., 2023) demonstrated that single policy coverage is sufficient for learning the optimal policy. However, we prove that this condition no longer holds for PbMARL by providing a counterexample. Instead, we introduce an algorithm that operates under unilateral coverage, a condition derived from offline MARL (Cui and Du, 2022a; Zhong et al., 2022). Specifically, this condition requires the dataset to cover all unilateral deviations from a Nash equilibrium policy. For further details, see Section 4.
Figure 1: The overall pipeline of offline PbMARL. $\mathcal{D}$ is the preference dataset where $\tau_{i},\tau_{i}^{\prime}$ are trajectories and $\mathbf{y}_{i}\in\{1,-1\}^{m}$ indicates which trajectory is preferred by each agent. $r_{\phi}$ is the learned reward. $\pi_{b}$ is the learned reference policy using imitation learning.
@@
As a foundational exploration into PbMARL research, we focus on employing the simplest learning framework, incorporating only the essential techniques necessary to ensure the approach’s feasibility. The framework consists of three key components: 1) leveraging the preference dataset to learn a reward function, 2) mitigating extrapolation errors with pessimism, and 3) determining the final policy. Figure 1 provides an overview of the process.
However, additional algorithmic techniques are required to identify a robust policy, even when the dataset demonstrates good coverage according to our theoretical insights.
@@
  • Dataset Distribution-Based Pessimism. To mitigate the extrapolation error in offline RL, we add an extra reward term based on the density of a certain state-action pair in the dataset to implement pessimism. In our approach, an imitation learning agent is trained to model the density function. The final policy is then trained using a DQN-based Value Decomposition Network (VDN) (Mnih et al., 2013; Sunehag et al., 2017). Our ablation study demonstrates the critical role of appropriately tuning the reward coefficient to ensure training stability and performance (see Table 4).
@@
Our experiments, following the pipeline described above, confirm the theoretical necessity of unilateral coverage. We performed extensive ablation studies across three Multi-Agent Particle Environment (MPE) scenarios—Spread-v3, Tag-v3, and Reference-v3 (Mordatch and Abbeel, 2017)—as well as the popular Overcooked environment (Carroll et al., 2020). These studies focused on the hyperparameter selection for the reward regularization coefficient $\alpha$, pessimism coefficient $\beta$, and dataset diversity. The empirical results (Table 2) demonstrate that: 1) augmenting expert demonstrations with trivial trajectories significantly improves performance, 2) unilateral datasets are advantageous, and 3) dataset diversity contributes to lower variance. Our ablation experiments underscore the effectiveness of the proposed algorithmic techniques. Additionally, we introduced a principled standardization technique that efficiently tunes hyperparameters across all environments and datasets.
@@ 2 Related Works
RLHF, or preference-based RL (PbRL), plays a pivotal role in alignment with various tasks such as video games (Warnell et al., 2018; Brown et al., 2019), robotics (Jain et al., 2013; Kupcsik et al., 2016; Christiano et al., 2023; Shin et al., 2023), image augmentation (Metcalf et al., 2024), and large language models (Ziegler et al., 2020; Wu et al., 2021; Nakano et al., 2022; Menick et al., 2022; Stiennon et al., 2022; Bai et al., 2022; Glaese et al., 2022; Ganguli et al., 2022; Ouyang et al., 2022). Additionally, a body of work focuses on the reward models behind preference data (Sadigh et al., 2017; Bıyık and Sadigh, 2018; Gao et al., 2022; Hejna and Sadigh, 2023). Recent works like VIPO (Cen et al., 2024) incorporates uncertainty-aware regularization into the reward model, while (Liu et al., 2024) address over-optimization using adversarial regularization. Direct preference optimization (DPO, Rafailov et al. (2023)) and its variants (Azar et al., 2023; Rafailov et al., 2024) approach RLHF without directly handling the reward model. Theoretical studies have also explored guarantees, such as sample complexity and regret, and the limitations of certain RLHF algorithms (Novoseller et al., 2020; Xu et al., 2020; Pacchiano et al., 2023; Chen et al., 2022; Razin et al., 2023; Zhu et al., 2024a; Wang et al., 2023c; Xiong et al., 2024; Zhu et al., 2024b).
@@
Offline RL (Lange et al., 2012; Levine et al., 2020) has achieved success in a wide range of real-world applications, including robotics (Pinto and Gupta, 2015; Levine et al., 2016; Chebotar et al., 2021; Kumar et al., 2023), healthcare (Raghu et al., 2017; Wang et al., 2018), and autonomous driving (Shi et al., 2021; Lee et al., 2024). Key algorithms such as Behavior Cloning, BRAC (Wu et al., 2019), BEAR (Kumar et al., 2019), and CQL (Kumar et al., 2020; Lyu et al., 2024) have driven these successes. Theoretical research on offline RL has primarily focused on sample complexity under various dataset coverage assumptions Le et al. (2019); Chen and Jiang (2019); Yin et al. (2020); Rashidinejad et al. (2023); Yin et al. (2021, 2022); Shi et al. (2022); Nguyen-Tang et al. (2022); Xie et al. (2022); Xiong et al. (2023b); Li et al. (2024); Xie et al. (2023); Mete et al. (2021).
@@
Many real-world scenarios are naturally modeled as multi-agent environments, whether cooperative or competitive. As a result, MARL has gained popularity in video games (Tian et al., 2017; Vinyals et al., 2017; Silver et al., 2017; Vinyals et al., 2019), network design (Shamsoshoara et al., 2018; Kaur and Kumar, 2020), energy sharing (Prasad and Dusparic, 2018), and autonomous driving (Palanisamy, 2019; Yu et al., 2020; Zhou et al., 2022). Prominent algorithms in MARL include IQL (Tan, 2003), MADDPG (Lowe et al., 2020), COMA (Foerster et al., 2017), MAPPO (Yu et al., 2022), VDN (Sunehag et al., 2017), and QMIX (Rashid et al., 2018). Theoretical research has made great process in reducing the sample complexity (Wang et al., 2023b; Xiong et al., 2023a).
@@
Offline MARL is a practical solution for handling sophisticated multi-agent environments. Empirically, to address issues related to out-of-distribution actions and complex reward functions, previous works have developed algorithms such as MABCQ (Jiang and Lu, 2023), ICQ-MA (Yang et al., 2021), OMAR (Pan et al., 2022), and OMIGA (Wang et al., 2023a), which incorporate regularization or constraints on these actions and functions. MOMA-PPO (Barde et al., 2024) is a model-based approach to offline MARL that generates synthetic interaction data from offline datasets. Tseng et al. (2022) combines knowledge distillation with multi-agent decision transformers (Meng et al., 2022) for offline MARL. Theoretical understanding of offline MARL, particularly in the context of Markov games, has been advanced by works that provide sample complexity guarantees for learning equilibria Sidford et al. (2019); Cui and Yang (2020); Zhang et al. (2023a, 2020); Abe and Kaneko (2020); Cui and Du (2022a, b); Zhang et al. (2023b); Blanchet et al. (2023); Shi et al. (2023); Zhong et al. (2022).
@@ 4 Dataset Coverage Theory for MARLHF
In this section, we study the dataset coverage assumptions for offline MARLHF. For offline single-agent RLHF, Zhu et al. (2023); Zhan et al. (2023) show that single policy coverage is sufficient for learning the optimal policy. However, we prove that this assumption is insufficient in the multi-agent setting by constructing an counterexample. In addition, we prove that unilateral policy coverage is adequate for learning the Nash equilibrium.
@@ 5 Algorithmic Techniques for Practical Performance
In Section 4, we provided a theoretical characterization of the dataset requirements for PbMARL. However, the algorithm used in Theorem 2 is not computationally efficient. In this section, we propose a practical algorithm for PbMARL and validate our theoretical findings through experiments.
@@ 5.1 High-level Methodology
Our PbMARL pipeline consists of two phases: In the first step, we train a reward prediction model $\phi$ and approximate the behavior policy $\pi_{b}$ using imitation learning; in the second step, we then apply an MARL algorithm to maximize a combination of the KL-divergence-based reward and standardized predicted reward $r_{\phi}$, ultimately deriving the final policy $\pi_{\mathbf{w}}$.
Step 1: Reward Training and Dataset Modeling. Given the preference signals of trajectories, we use neural networks to predict step-wise rewards $r_{\phi}(s_{h},a_{h})$ for each agent, minimizing the loss defined in (1). The objective is to map $(s,a_{i})$-pairs to reward values such that the team returns align with the preference signals. At the same time, in order to utilize distribution-based penalty term $\log{\pi_{b}(s,a)}$ to cope with the extrapolation error in offline learning, an imitation learner is trained over the entire dataset to model the behavior policy $\pi_{b}$.
Step 2: Offline MARL. Although in this work, VDN is chosen as the MARL oracle, it should be noted that other MARL architectures are also applicable. With the reward model $r_{\phi}$ and the approximated dataset distribution learned in Step 1, we are now able to construct a virtual step-wise reward for each agent. The agents are then trained to maximize the target defined in (3).
@@ 5.2 Reward Regularization
5.3 Dataset Distribution-Based Pessimism
There are various methods to mitigate the over-extrapolation errors in offline RL (Peng et al., 2019; Nair et al., 2021), including conservative loss over the Q-function (Kumar et al., 2020) and directly restricting the learned policy actions to those within within the dataset (Fujimoto et al., 2019). We add a per-step dataset-based penalty term, $\log{\pi_{b}(s,\mathbf{a})}$, as pessimism towards less explored states. Imitation learning is utilized to estimate the behavior policy $\pi_{b}$ from the dataset distribution. To stabilize training, we standardize predicted reward $r_{\phi}$ over $\mathcal{D}$ before combining it with the penalty term to make them comparable:
$\text{objective}(\mathbf{w})=\mathbb{E}_{\tau\sim\pi_{\mathbf{w}}}\left[\sum_{h=1}^{H}r_{\text{std}}(s_{h},\mathbf{a}_{h},\phi)+\beta\log{\pi_{b}(s_{h},\mathbf{a}_{h})}\right],$ (3)
[…]
@@ 5.3 Dataset Distribution-Based Pessimism
Intuitively, the penalty term $\log{\pi_{b}(s_{h},\mathbf{a}_{h})}$ discourages the agents from deviating from the most preferred actions in the dataset. The effectiveness of this method is validated in the ablation study (cf. Section 6.3).
@@ 6 Experiments
We design a series of experiments to validate our theories and methods in common general-sum games. Specifically, we first use online RL algorithms to train expert agents, and take intermediate checkpoints as rookie agents. Then, we use these agents to collect datasets and use the Bradley-Terry model over standardized returns to simulate human preference. Experiments are carried out to verify the efficiency of our approach with unilateral policy dataset coverage (in Theorem 2) while single policy coverage is insufficient (stated in Theorem 1). We also design ablation studies to showcase the importance of our methods, particularly focusing on reward regularization and dataset distribution-based pessimism.